Bob Grey invests $22,000, some at 7% and some at 6% annual interest. If he recieves an annual return of $1,420, how much should be invested at eact rate?
Answer by rfer(16322)

0.07x+0.06(22000-x)=1420
0.07x+1320-0.06x=1420
0.01x=100
x=10,000
22000-x=12,000